# Machine Learning for the Physical Sciences

### v1

@@ 2 Dataset
Our dataset aligns heterogeneous data from a diverse set of sources, encompassing multiple modalities, cadences, and start/end dates. Data sources include global ionospheric maps (GIMs) hernandez2009igs , both with sparse and dense measurements, Solar Dynamics Observatory (SDO) extreme ultraviolet flux embeddings, solar driver data, and geomagnetic driver data. Our data product aligns these data sources in time and incorporates relevant orbital mechanics and quasi-dipole features. The aligned data product is visualized in Figure 1, and is publicly available as a Google Cloud bucket  [^1] https://console.cloud.google.com/storage/browser/ionosphere_data_public.
@@ 4 Baseline Results
Our data product is available publicly as a Google Cloud bucket, along with the codebase used to align, process, and split data based on geomagnetic storm events  [^3] https://console.cloud.google.com/storage/browser/ionosphere_data_public. Our codebase includes PyTorch datasets that prepare training data according to user-specified start and end ranges, dataset-specific normalization schemes, and example PyTorch model training code.

### v2

@@ 2 Dataset
Our dataset aligns heterogeneous data from a diverse set of sources, encompassing multiple modalities, cadences, and start/end dates. Data sources include global ionospheric maps (GIMs) hernandez2009igs , both with sparse and dense measurements, Solar Dynamics Observatory (SDO) extreme ultraviolet flux embeddings, solar driver data, and geomagnetic driver data. Our data product aligns these data sources in time and incorporates relevant orbital mechanics and quasi-dipole features. The aligned data product is visualized in Figure 1, and is publicly available on AWS S3  [^1] https://zenodo.org/records/18343833 and https://github.com/FrontierDevelopmentLab/2025-HL-Ionosphere-dataset.
@@ 4 Baseline Results
Our data product is available publicly on AWS S3, along with the codebase used to align, process, and split data based on geomagnetic storm events  [^3] https://zenodo.org/records/18343833 and https://github.com/FrontierDevelopmentLab/2025-HL-Ionosphere-dataset. Our codebase includes PyTorch datasets that prepare training data according to user-specified start and end ranges, dataset-specific normalization schemes, and example PyTorch model training code.